# Data-Driven Computation of Polytopic Invariant Sets for Noisy Nonlinear Systems

### v1

@@ 5. Simulation Results
where the parameters are set to $\alpha=0.01$ and $\mu=0.9$. These parameters are considered to be unknown throught the whole simulations. The constraints on the input and state are given by $U=\{u\in\mathbb{R}:|u|\leq 2\}$ and $X=\{x\in\mathbb{R}^{2}:\|x\|_{\infty}\leq 4\}$, respectively. These systems’ parameters are unknown to us. The uncertainty set is bounded by $W=\{w\in\mathbb{R}^{n}:\|w\|_{\infty}\leq 0.4\}$, which the upper bound is known.
@@
The numerical simulations presented in this section were carried out in Python and Matlab environment on a computer equipped with an Apple M4 processor and 32 GB of RAM.
@@
To construct the consistency set defined in (5), we first generate a dataset by sampling trajectories from the system (16), collecting a total of 30 state-input pairs. Once the consistency set is constructed, its vertices are calculated using the con2vert function in MATLAB. For the given dataset size and noise bound, this computation yields 667 vertices and takes approximately 1.3 seconds. Now, we need to shape the function $F$ as mentioned in (9). To do this, one of the possible choices for the decomposition can be the following:
$g_{1}(x_{k},u_{k})$ $=a_{1}x_{1,k}+a_{2}x_{2,k}+b_{1}u_{k}+\frac{a_{3}}{4}(x_{1,k}+u_{k})^{2}$
$h_{1}(x_{k},u_{k})$ $=\frac{a_{3}}{4}(x_{1,k}-u_{k})^{2}$
$g_{2}(x_{k},u_{k})$ $=a_{4}x_{1,k}+a_{5}x_{2,k}+b_{2}u_{k}+\frac{a_{6}}{4}(x_{2,k}-u_{k})^{2}$
$h_{2}(x_{k},u_{k})$ $=\frac{a_{6}}{4}(x_{2,k}+u_{k})^{2}$
where $a_{1},a_{2},\cdots,a_{6},b_{1},b_{2}$ are unknown parameters. We replace these parameters with the calculated vertices of consistency set. By substituing the vertices on the decomposition and shape the function $F$, we are able to straightly use Algorithm 1 to find the largest polytope within $X$.
@@
Figure 1. Data-driven computation of largest contractive set (purple solid line) within the state constraint set.
@@
Figure 1 illustrates the computational results of Algorithm 1. The outer box represents the state constraint set, defined by $\|x\|_{\infty}\leq 4$. The small, dashed-line ellipse at the center is the initial candidate polytope, $\Omega$, that provides the initial shape for the algorithm. The large, purple shaded region is the final robustly $\lambda$-contractive set computed by our data-driven method. This set is the result of scaling the initial polytope $\Omega$ by the factor $\alpha=11.264$, which was determined by the algorithm to be the largest possible scaling that guarantees robust invariance for all systems compatible with available information.
While the computed set shown in Figure 1 is large, it is not necessarily the maximal invariant set. The conservatism of this result can be further reduced by applying an iterative enlarging method, similar to the one proposed by the authors as discussed in Remark 6.
Based on the enlarging method, we have the result is shown in Fig 2.
Figure 2. Data-driven computation of largest contarctive set using randomized enlarging method.
It can be seen in Fig. 2 that after adding 20 random points within the state constraints set, less conservative $\lambda$-contractive set is achieved.
6. Conclusion
In this paper, we have presented a novel data-driven framework for computing robust polytopic convex $\lambda$-contractive sets for noisy nonlinear systems. The proposed approach successfully addresses the challenge of synthesizing provably invariant sets when a mathematical model of the system is unavailable, relying only on a finite set of noisy measurements. By constructing a polytopic consistency set to capture all possible system compatible with available infromation, we adapted a sufficient condition from the model-based literature, originally based on DC functions, to this data-driven context.
The effectiveness of the proposed algorithm was demonstrated through a numerical example, successfully computing the largest possible contractive set in state constraint set. Future research will focus on two key areas. First, a critical next step is to develop methods that mitigate the computational complexity, which scales with the number of vertices in the consistency and initial contractive set. Second, future work should address the conservatism of the approach. As the DC decomposition of the system dynamics is not unique, developing a systematic method to find an optimal decomposition that yields the largest possible $\lambda$-contractive set is a promising direction for enhancing the framework.

### v2

@@ V Simulation Results
where the parameters are $\tau=0.01$ and $\mu=0.9$. The parameters $\{a_{1},a_{2},\cdots,b_{1},b_{2}\}$ are assumed unknown throughout the simulation. The constraints on the input and state are given by $U=\{u\in\mathbb{R}:|u|\leq 2\}$ and $X=\{x\in\mathbb{R}^{2}:\|x\|_{\infty}\leq 4\}$, respectively. The uncertainty set for the continuous system is bounded by $W=\{w\in\mathbb{R}^{n}:\|w\|_{\infty}\leq 0.4\}$, for which the upper bound is known. The data dictionary, $\mathcal{D}=\{(x_{k},u_{k},x_{k+1})\}_{k=0}^{k=30}$, is generated using (12).
@@
Simulation Environment: The numerical simulations presented were carried out in Python/Matlab environment on an Apple M4 processor and 32 GB of RAM.
@@
To construct the consistency set defined in (5), we first generate a dataset by sampling trajectories from the system (12), collecting a total of 30 state-input pairs. Once the consistency set is constructed, its vertices are calculated using the con2vert function in MATLAB. For the given dataset size and noise bound, this computation yields $n_{q}=667$ vertices and takes approximately 1.3 seconds. Now, we need to construct the function $F$ as mentioned in (9). To do this, one possible choice for the decomposition can be the following:
$g_{1}(x_{k},u_{k})$ $=a_{1}x_{1,k}+a_{2}x_{2,k}+b_{1}u_{k}$ (14)
$\quad+\frac{|a_{3}|}{4}\big(x_{1,k}+\text{sgn}(a_{3})u_{k}\big)^{2},$
$h_{1}(x_{k},u_{k})$ $=\frac{|a_{3}|}{4}\big(x_{1,k}-\text{sgn}(a_{3})u_{k}\big)^{2},$
$g_{2}(x_{k},u_{k})$ $=a_{4}x_{1,k}+a_{5}x_{2,k}+b_{2}u_{k}$
$\quad+\frac{|a_{6}|}{4}\big(x_{2,k}-\text{sgn}(a_{6})u_{k}\big)^{2},$
$h_{2}(x_{k},u_{k})$ $=\frac{|a_{6}|}{4}\big(x_{2,k}+\text{sgn}(a_{6})u_{k}\big)^{2},$
Recall that the true parameter vector $\theta=[a_{1},a_{2},a_{3},a_{4},a_{5},a_{6},b_{1},b_{2}]^{T}$ is unknown. Therefore, to evaluate the robust invariance condition in Theorem 1, the parameters in (14) are replaced by the corresponding elements of each consistency set vertex $\theta^{*,s}$ for $s\in\mathbb{N}_{n_{q}}$. By substituting these vertex parameters into the decomposition and constructing the parameterized function $F(\cdot,\cdot,\cdot;\theta^{*,s})$ for every calculated vertex, we are able to use Algorithm 1 to find the largest possible polytope within $X$.
@@
Figure 1: Data-driven computation of largest contractive set (purple solid line) within the state constraint set.
@@
Figure 1 illustrates the computational results of Algorithm 1. The outer box represents the state constraint set, defined by $\|x\|_{\infty}\leq 4$. The small, dashed-line ellipse at the center is the initial candidate polytope, $\Omega_{0}$, that provides the initial shape for the algorithm. The large, purple shaded region is the final robustly $\lambda_{\omega}$-contractive set computed by our data-driven method. This set is the result of scaling the initial polytope $\Omega_{0}$ by the factor $\alpha=8.57$, which was determined by the algorithm to be the largest possible scaling that guarantees robust invariance for all systems compatible with available information. Although the data-driven $\lambda_{\omega}$-contractive set computed in Fig. 1 is notably large, it remains a conservative inner approximation of the true maximal invariant set. This conservatism can be systematically reduced by employing the iterative enlarging technique detailed in Remark 7. The green enlarged polytope presents the expanded set obtained through this procedure. Specifically, the introduction of 20 random candidate points within the allowable state constraints allows the algorithm to augment the original polytope, resulting in a demonstrably less conservative $\lambda_{\omega}$-contractive set.
VI Conclusion
This paper introduced a novel data-driven framework for the approximation of robust polytopic $\lambda_{\omega}$-contractive sets in uncertain nonlinear systems. The proposed methodology overcomes the limitations of traditional model-based techniques by synthesizing provably invariant sets directly from a finite sequence of noisy state-input measurements. By leveraging the properties of DC functions, we successfully extended established model-based sufficient conditions for set invariance to the data-driven domain. A numerical case study validated the tractability and efficacy of the proposed optimization-based algorithm, demonstrating its potential for robust control applications.
VII Appendix
Proof of Theorem 1.
Under Assumption 4, the consistency set is a compact polytope. This implies that for any $\theta$ within (5), we can express $\theta$ as a convex combination of the set’s vertices, such that $\theta=\sum_{s=1}^{n_{q}}\beta_{s}\theta^{*,s}$, with coefficients $\beta_{s}\geq 0$ and $\sum_{s=1}^{n_{q}}\beta_{s}=1$. Therefore, each vertex of the consistency set is interpreted as a ”worst-case” realization of the unknown system parameters. Now, the dynamics associated with the parameter vector $\theta$ within the consistency set can be written as follows:
$f(x_{k},u_{k};\theta)=$ $\sum_{s=1}^{n_{q}}\beta_{s}g(x_{k},u_{k};\theta^{*,s})-\sum_{s=1}^{n_{q}}\beta_{s}h(x_{k},u_{k};\theta^{*,s}),$ (15)
where $g,h$ are the decomposed convex vector functions of the system dynamics associated with the $\theta^{*,s}$ vertex. Let (15) be the decomposition for system corresponding to $\theta^{*,s}$ vertex in consistency set and the definition of $F(x,u,c;\theta^{*,s})$ proposed in (9), we can write:
$F(x,u,c;\theta)=$ $\sum_{s=1}^{n_{q}}\beta_{s}F(x,u,c;\theta^{*,s}).$ (16)
Based on Property (1), the function $F$ is convex. Considering the initial polytope $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\mathbf{1}\}\subseteq X$, we reach at the following relationship:
$F(v^{p},u^{p}_{v},H_{i}^{T};\theta)=\sum_{s=1}^{n_{q}}\beta_{s}F(v^{p},u^{p}_{v},H_{i}^{T};\theta^{*,s}),$ (17)
where $v^{p}$ denotes the vertex of $\Omega$, and $u^{p}_{v}$ represents the input of the system at $v^{p}$ vertex. We can now adapt the sufficient condition for $\lambda_{\omega}$-contractiveness, originally proposed for known nonlinear systems in [8] (Property 3). It can be said that each vertex $\theta^{*,s}$ is associated with a worst case system parameters in consistency set and if for all $\beta_{s}$ in (16), the condition (3) holds, the set $\Omega$ will be $\lambda_{\omega}$-contractive for all systems compatible with the available information.
It remains to show that there exists a control law $u$ such that for any initial state $x_{0}\in\Omega$, the trajectory of the system (1) subject to process noise satisfies $x_{k}\in\lambda_{\omega}^{k}\Omega$ for all $k\geq 0$. To prove the existence of an admissible control law for any initial state $x\in\Omega$, we explicitly construct it as follows: Since $\Omega$ is a polytope, any $x$ can be expressed as a convex combination of its vertices $\{v^{1},\dots,v^{n_{v}}\}$:
$x=\sum_{p=1}^{n_{v}}\zeta_{p}v^{p},\quad\text{with }\zeta_{p}\geq 0,\sum_{p=1}^{n_{v}}\zeta_{p}=1,$
Let $u^{p}_{v}$ be the control actions at the vertices that satisfy the condition in Theorem 1. We define the control law $u\in U$ using the same convex combination coefficients $\zeta_{p}$:
$u=\sum_{p=1}^{n_{v}}\zeta_{p}u^{p}_{v},$
Next, we must show that for any $x_{0}\in\Omega$ and any disturbance $w_{0}\in W$, the successor state $x_{k+1}=f(x_{k},u_{k};\theta)+w_{k}$ is contained in $\lambda_{\omega}\Omega$ for all $\theta$ in the consistency set $\mathcal{P}_{1}$.
From the properties of the support function and the function $F$, we have the following upper bound for all $\theta$, and $\forall i\in\mathbb{N}_{n_{h}}$:
$H_{i}(f(x_{k},u_{k};\theta)+w_{k})\leq F(x_{k},u_{k},H_{i}^{T};\theta)+\phi_{W}(H_{i}^{T}),$ (18)
As established, the function $F(x,u,c;\theta)$ is convex in the pair $(x,u)$. Using (17) and as shown in the proof of Theorem 1, the condition $F(v^{p},u^{p}_{v},H_{i}^{T};\theta)\leq\lambda_{\omega}-\phi_{W}(H_{i}^{T})$ holds for all $\theta$ in consistency set. We can write:
$H_{i}x_{k+1}$ $\leq\left(\sum_{p=1}^{n_{v}}\zeta_{p}F(v^{p},u^{p}_{v},H_{i}^{T};\theta)\right)+\phi_{W}(H_{i}^{T})$
$\leq\left(\sum_{p=1}^{n_{v}}\zeta_{p}(\lambda_{\omega}-\phi_{W}(H_{i}^{T}))\right)+\phi_{W}(H_{i}^{T})$
$=(\lambda_{\omega}-\phi_{W}(H_{i}^{T}))\sum_{p=1}^{n_{v}}\zeta_{p}+\phi_{W}(H_{i}^{T})$
$=\lambda_{\omega}-\phi_{W}(H_{i}^{T})+\phi_{W}(H_{i}^{T})=\lambda_{\omega},$
Since $H_{i}x_{k+1}\leq\lambda_{\omega}$ for all facets $i\in\mathbb{N}_{n_{h}}$ and vertices of consistency set, we have shown that $x_{k+1}\in\lambda_{\omega}\Omega$. ∎